### v1

# Online Density-Based Clustering for Real-Time Narrative Evolution Monitoring

Abstract
Automated narrative intelligence systems for social media monitoring face significant scalability challenges when processing continuous data streams using traditional batch clustering algorithms. We investigate the replacement of HDBSCAN (offline clustering) with online (streaming/incremental) clustering methods in a production narrative report generation pipeline. The proposed system employs a three-stage architecture (data collection, modeling, dashboard generation) that processes thousands of multilingual social media documents daily. While HDBSCAN excels at discovering hierarchical density-based clusters and handling noise, its batch-only nature necessitates complete retraining for each time window, resulting in memory constraints, computational inefficiency, and inability to adapt to evolving narratives in real-time. This work evaluates a bunch of online clustering algorithms across dimensions of cluster quality preservation, computational efficiency, memory footprint, and integration compatibility with existing workflows. We propose evaluation criteria that balance traditional clustering metrics (Silhouette Coefficient, Davies-Bouldin Index) with narrative metrics (narrative distinctness, contingency and variance). Our methodology includes sliding-window simulations on historical datasets from Ukraine information space, enabling comparative analysis of algorithmic trade-offs in realistic operational contexts. This research addresses a critical gap between batch-oriented topic modeling frameworks and the streaming nature of social media monitoring, with implications for computational social science, crisis informatics, and narrative surveillance systems.
Online Density-Based Clustering for Real-Time Narrative Evolution Monitoring
Ostap Vykhopen, Viktoria Skorik, Maxim Tereschenko, Veronika Solopova ostap.vykhopen@mantisanalytics.com
1 Introduction
Automated narrative intelligence systems have become essential tools for monitoring information ecosystems, particularly in contexts requiring rapid detection of emerging narratives, misinformation campaigns, and geopolitical discourse patterns. Organizations ranging from defense agencies to public health authorities rely on these systems to process high-volume social media streams and generate actionable intelligence reports. However, the computational architecture of such systems often reflects fundamental tensions between the batch-oriented nature of modern machine learning frameworks and the inherently streaming character of social media data. This tension becomes particularly acute when examining HDBSCAN McInnes et al. (2017), a clustering algorithm widely adopted for narrative discovery, especially popularised by BerTopic library Scarpino et al. (2022); Eklund et al. (2023).
Despite its theoretical advantages, including the ability to handle varying density clusters, detect noise, and provide hierarchical structure, HDBSCAN’s strict batch processing requirement creates significant operational bottlenecks in production environments: (i) Memory constraints. Each clustering operation requires loading complete document embeddings (N × 100 dimensional UMAP reduced vectors) into memory, with O(NlogN) average case complexity for minimum spanning tree construction; (ii) computational inefficiency. For example, daily monitoring require complete re-clustering of new data each day, discarding previously learned cluster structures even as narrative topics evolve gradually; (iii) Parameter sensitivity. Production logs reveal frequent cluster size parameter tuning, indicating difficulty in maintaining stable cluster granularity across varying data volumes; (iv) inability to handle incremental updates. When new documents arrive, the entire clustering process must restart, preventing real-time cluster assignment or progressive refinement of narrative understanding; (v) scalability limitations. As historical data accumulates, the batch processing approach becomes increasingly untenable.
These limitations motivate our central research question: Can online clustering algorithms effectively replace HDBSCAN in narrative monitoring pipelines while preserving cluster quality, maintaining system integration compatibility, and improving computational efficiency? Specifically, we test clustering algorithms that not only support online mode but also automatically determine the number of clusters without requiring predefined cluster counts (as required by batch K-means) and demonstrate robustness to noise inherent in social media data streams. Furthermore, the ideal online clustering approach should achieve significant efficiency improvements, where the memory footprint of the online algorithm is substantially smaller than HDBSCAN’s batch requirements, and the time required for incremental updates is considerably faster than HDBSCAN’s complete re-clustering operations.
Our contributions are threefold:
  1. 1.
    We introduce a realistic experimental framework for evaluating clustering under streaming conditions, including temporal pretraining, incremental daily updates, and sliding-window simulations on high-volume multilingual social media data.
  2. 2.
    We conduct a systematic comparison of online density-based algorithms (DBSTREAM, DenStream) against the HDBSCAN baseline within a production-style pipeline, using a joint evaluation scheme that combines traditional clustering metrics (Silhouette, Davies–Bouldin) with narrative-specific metrics (distinctness, contingency, variance)
  3. 3.
    We identify implementation-level scalability and correctness issues in widely used reference libraries and show how these affect real-world performance. Based on empirical and operational analysis, we position DenStream as the most promising online alternative and outline the architectural and downstream adaptations required for deployment in streaming narrative monitoring systems.
2 Theoretical Background and Related work
2.1 Density-Based Clustering
Density-based clustering algorithms, pioneered by Density-Based Spatial Clustering of Applications with Noise (DBSCAN) from Ester et al. (1996), have become foundational for discovering arbitrary-shaped clusters in spatial data while robustly handling noise. Unlike partition-based methods (k-means) or hierarchical agglomerative clustering, DBSCAN defines clusters as dense regions separated by sparse areas, requiring only two parameters: $\epsilon$ (neighborhood radius) and min_samples (minimum points to form a dense region).
HDBSCAN extends DBSCAN by eliminating the need for $\epsilon$ through hierarchical cluster extraction based on cluster stability. The algorithm constructs a minimum spanning tree over mutual reachability distances, then extracts a flat clustering by selecting stable branches of the hierarchical tree. This approach provides several advantages: (1) automatic determination of optimal cluster granularity, (2) varying density cluster detection, (3) hierarchical cluster structure, and (4) probabilistic soft cluster assignments via outlier scores.
Despite these strengths, traditional implementations of DBSCAN and HDBSCAN require access to the complete dataset to construct their distance-based structures and compute cluster hierarchies. This computational model assumes static data and the ability to hold all points in memory simultaneously, limiting applicability to streaming scenarios where data arrives continuously and retrospective reprocessing of the entire history is impractical. Furthermore, narrative monitoring applications, where topics emerge, evolve, and fade over time, demand algorithms that can discover clusters without predefining their number, adapting dynamically to the natural structure present in incoming document streams rather than imposing predetermined partitions.
2.2 Online and Streaming Clustering Algorithms
The streaming clustering literature addresses scenarios where data arrives continuously and memory constraints prevent storing the entire dataset.
CluStream Chan et al. (1982); Caussinus et al. (1982) is one of the earliest and most influential data stream clustering frameworks, introducing a two-phase approach consisting of online micro-clustering and offline macro-clustering. In the online phase, it maintains temporal micro-clusters using summary statistics and timestamps, while the offline phase applies traditional clustering algorithms (e.g., k-means) over these summaries to generate final clusters at different time horizons. While CluStream supports temporal analysis and evolving streams, it requires the number of clusters to be specified in advance and is primarily suited for spherical cluster structures.
DBSTREAM Hahsler and Bolanos (2016) is a density-based data stream clustering algorithm that extends micro-cluster approaches by maintaining a shared density graph capturing the actual density between adjacent micro-clusters during the online summarization. At reclustering time, it uses this shared density information to connect micro-clusters into final clusters more accurately than methods relying only on distances or assumptions about internal distributions.
DenStream Cao et al. (2006) is another density-based stream clustering maintaining micro-clusters (core-micro-clusters and outlier-micro-clusters) with temporal decay weights. Its crucial advantage is that it addresses concept drift by gradually forgetting old data and it supports arbitrary cluster shapes like DBSCAN. However, this introduces additional complexity in managing micro-cluster lifecycle and decay parameters.
Recent neural approaches, including deep clustering networks and variational autoencoders with clustering objectives, offer representation learning combined with clustering, yet they typically require large training datasets and provide limited interpretability, which is essential for narrative monitoring applications.
2.3 Streaming Topic Modeling
Traditional topic modeling methods such as LDA and NMF have been extended to streaming and online contexts using incremental variational inference and related techniques, and researchers have also developed enhanced versions that combine topic models with semantic representations Das (2025). However, the canonical implementations of LDA and NMF still operate on bag‑of‑words or weighted document‑term matrices and do not inherently capture deep contextual similarity like transformer‑based dense embeddings do, motivating hybrid models that merge topic modeling with embedding spaces for richer semantic coherence Yadav et al. (2025).
Recent advances in neural embedding models have enabled a fundamentally different approach to streaming topic modeling that leverages the rich semantic representations of transformer-based architectures Angelov (2020); Bianchi et al. (2021). Rather than relying on term-frequency statistics, our approach builds on a pipeline that transforms documents into dense embeddings, applies dimensionality reduction to preserve local semantic structure in a lower-dimensional space, performs clustering to identify topically coherent groups, and finally employs large language models to generate interpretable topic labels and summaries. This embedding-first paradigm naturally captures contextual similarity and semantic nuance while remaining compatible with streaming scenarios through incremental clustering updates. By decoupling representation learning from cluster assignment, this framework addresses key limitations of traditional topic models, such as their inability to capture deep semantic relationships, while maintaining the flexibility needed for online, continuously arriving data streams.
3 Methodology & Experimental setup
Figure 1: Four stage pipeline for topic and narrative discovery.
3.1 System Architecture
Our narrative intelligence system implements a four-stage pipeline for automated topic discovery and monitoring in social media streams.
Data Collection and Preprocessing.
Raw documents are collected from Ukrainian information space sources, including news portals, Telegram channels, and user comments. Text undergoes standard preprocessing including normalization, and language detection filtering.
Modeling Pipeline.
The core topic modeling pipeline consists of four sequential stages, and is illustrated in Figure 1:
  1. 1.
    Embedding Generation: For all experiments, the documents are transformed into dense semantic representations using multilingual MiniLM111sentence-transformers/all-MiniLM-L6-v2 model from sentence transformers library.
  2. 2.
    Dimensionality Reduction: Document embeddings undergo dimensionality reduction (typically via UMAP), preserving local neighborhood structure while reducing computational complexity for downstream clustering
  3. 3.
    Clustering: Documents are grouped into topically coherent clusters based on their reduced embeddings. This is the component under investigation in our experiments
  4. 4.
    LLM-based Topic Labeling and Enrichment: to isolate the influence of clustering algorithms, a proprietary LLM, with stable parameters and prompt was used to generates human-interpretable topic labels and summaries for each discovered cluster in all experimental settings.
The final component provides operational intelligence visualizations and reporting interfaces for analysts This pipeline architecture decouples representation learning from cluster discovery, enabling modular replacement of clustering algorithms while maintaining embedding quality and downstream interpretation capabilities.
3.2 Experiment Design
Dataset
We evaluate online clustering algorithms using data from the Ukrainian information space, including news portals, Telegram channels, and user comments. The dataset [^2] The data can be shared upon request to vetted academics for reproducibility purposes. exhibits typical characteristics of narrative monitoring applications: high volume ($\approx$17,000 documents per day), topical diversity reflecting ongoing geopolitical discourse, and substantial noise from low-quality or off-topic content.
Table 1: Clustering Performance Comparison. The table reports clustering quality metrics for the baseline HDBSCAN algorithm and two online clustering approaches evaluated on daily data from the Ukrainian information space (approximately $N\approx 17{,}000$ documents per day before cleaning and filtering).
Algorithm Silhouette DBI Distinct. Contingency Variance
HDBSCAN 0.592 0.550 0.352 0.299 0.286
DBSTREAM 0.327 1.220 0.266 0.630 0.459
DenStream 0.685 0.453 0.319 0.389 0.319
Table 2: Operational Metrics. The number of narrative candidates does not represent the actual number of narratives, but rather the number of clusters classified as narratives by the model. This distinction should be considered when interpreting the results.
Algorithm Train (s) Pred. (s) Clusters Narratives
HDBSCAN 13 1.3 1063 118
DBSTREAM 232 127 266 58
DenStream 3.56 1.7 303 53
(a) HDBSCAN
(b) DBSTREAM
(c) DenStream
Figure 2: Comparison of online clustering methods: (a) HDBSCAN, (b) DBSTREAM, and (c) DenStream. Each fingerprint represents a cluster projected onto a set of domain-specific keywords, forming a polygon that encodes its semantic profile. Sharp spikes indicate focused narratives, while broad shapes show multi-topic or hybrid narratives. Overlapping areas reveal shared or competing narratives, and may also indicate overlapping clustering. The charts above show the Semantic Fingerprints of three clusters, based on a set of highly relevant keywords in the current Ukrainian information space. Chart (A) (HDBSCAN) shows that the first cluster covers nearly all keywords, which may indicate contamination by a mix of different narratives. In contrast, DBSTREAM and DenStream display a clearer per-keyword distribution and better separation between clusters.
For the experiments, data are split to reflect realistic streaming conditions. A pretraining set of 68,917 documents collected over six days (01.05–01.11) is used to initialise the models, followed by a streaming training day comprising 11,466 documents from 01.12, on which online updates and clustering behaviour are evaluated.
Baseline System.
Our baseline employs HDBSCAN operating in strict global mode, where clustering is performed independently on each day’s data without reference to historical patterns. Specifically, for a target day d, HDBSCAN processes only the documents from day d, discovering clusters de novo without leveraging any structure learned from previous days. This approach represents the current production paradigm where each daily report requires complete re-clustering.
Online Algorithm Evaluation.
We evaluate three online clustering algorithms from the River library: DBSTREAM, DenStream, TextClust. Each online algorithm follows an incremental training protocol: (1) initialization on historical context (6 days of data preceding the target day), and (2) incremental updating and prediction on the day of interest. This approach simulates a realistic production deployment where the system has access to recent historical context before processing new daily data streams.
Evaluation Metrics.
We assess clustering quality using three standard metrics computed on the day of interest: (i) Silhouette Score [-1, 1], which measures cluster cohesion and separation, with higher values indicating better-defined clusters; (ii) Davies-Bouldin Index [0, $\infty$], measuring average similarity between clusters and their most similar cluster, with lower values indicating better separation; (iii) Narrative Distinctness Irani et al. (2025), which quantifies how distinct a narrative is by combining the average and minimum cosine distances between its centroid and all other narrative centroids and (iv) Narrative Contingency and (v) narrative variance Upravitelev et al. (2025), measuring the cosine distance between a centroid and the top-ranked embedding from the set and showing the variance of text embeddings relative to their centroid, respectively. $\sigma^{2}=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\|\mathbf{t}_{ij}-\mathbf{c}_{i}\|_{2}^{2}$
Additionally, we measure computational efficiency through training time, prediction time, and the number of discovered clusters. These operational metrics directly address the scalability concerns motivating our investigation.
4 Results
We demonstrate our results in Table 1 and 2, while the difference in semantic fingerprints of the top 3 clusters of each algorithm can be seen in Figure 2.
4.1 Cluster Quality Metrics
DenStream demonstrates superior cluster quality across conventional clustering metrics. It achieves the highest Silhouette Score (0.685), indicating well separated and cohesive clusters, representing improvement over the HDBSCAN baseline (0.592). Similarly, DenStream attains the lowest Davies-Bouldin Index (0.453) compared to the baseline (0.550), indicating better cluster separation with lower inter-cluster similarity. DBSTREAM exhibits poor performance on standard clustering quality metrics. Its Silhouette Score (0.327) is notably lower than the baseline (0.592), while its Davies-Bouldin Index (1.220) is more than twice the baseline value (0.550), indicating substantial inter-cluster confusion. This suggests DBSTREAM’s density based micro-cluster approach may be insufficiently granular for the high dimensional semantic space of document embeddings.
4.2 Narrative-Specific Metrics
The narrative specific metrics reveal a more nuanced picture of algorithm behavior in the context of topic discovery for intelligence applications. Narrative Distinctness measures the uniqueness and separability of discovered narrative themes. While the baseline achieves the highest score (0.352), DenStream (0.319) outperforms DBSTREAM (0.266), indicating that DenStream is more effective than DBSTREAM at identifying distinct topical narratives, though it does not surpass the baseline. The baseline HDBSCAN achieves the lowest Contingency (0.299) and Variance (0.286), suggesting stable, predictable cluster structures within each daily batch. In contrast, both online algorithms exhibit substantially higher values (DenStream: 0.389, 0.319; DBSTREAM: 0.630, 0.459), indicating greater variability in narrative structure across time. This increased contingency likely reflects the incremental nature of online algorithms, which adapt cluster structures as new documents arrive rather than performing global optimization on complete daily batches.
HDBSCAN produces three times more clusters than DenStream, even though the Number of Narrative Candidates differs between them by only a factor of two. This may indicate excessive granularity in HDBSCAN, manifested in the fragmentation of a single narrative into multiple distinct clusters. Conversely, DenStream may tend to merge several distinct narratives into a single cluster.
5 Discussion
Overall, our results indicate that DenStream provides the most robust clustering performance, achieving clear improvements over the HDBSCAN baseline in both conventional and narrative-specific metrics. Its higher Silhouette Score and lower Davies–Bouldin Index suggest well-separated and cohesive clusters, which is further reflected in its superior Narrative Distinctness, highlighting its ability to recover meaningful and distinct thematic structures. In contrast, DBSTREAM performs poorly across standard clustering metrics, likely due to limitations of its micro-cluster representation in high-dimensional semantic embedding spaces. While the baseline exhibits greater temporal stability, as indicated by lower Narrative Contingency and Variance, the increased variability observed for the online algorithms appears to be an inherent trade-off of incremental clustering. This adaptability may be advantageous in dynamic intelligence settings, where evolving narratives are expected rather than static topic structures.
However, during the implementation and experimental evaluation of clustering algorithms, several technical limitations were identified in commonly used reference implementations, substantially affecting performance and scalability, which should be taken into consideration when choosing a method. For DenStream, the predict_one method in the River ML library recomputes final clusters by executing a DBSCAN-like procedure over all micro-clusters on every prediction call. While this behavior is consistent with fully online streaming assumptions, it leads to prohibitively slow prediction when the model state is static. This limitation is acknowledged by the River maintainers, who note that ‘DenStream.predict_one‘ performs excessive recomputation for certain use cases. In offline evaluation settings, this repeated reclustering is unnecessary. Additionally, inconsistencies were observed in River’s DenstreamMicroCluster implementation, particularly in the maintenance of running weighted statistics, which can result in incorrect micro-cluster centers, radii, and weights.
For CluStream, performance issues arise during training rather than prediction. When micro-cluster merging is required, the algorithm performs quadratic-time pairwise distance computations between all micro-cluster centers, resulting in (O($M^{2}$)) complexity per merge event Mansalis et al. (2018). Combined with additional overhead from periodic k-means reclustering, this leads to rapidly increasing training time as the stream progresses. Due to these inherent scalability limitations, CluStream was fully rejected from further consideration in this study.
As a result of the identified performance and correctness issues, the River implementations were deemed unsuitable for large-scale experiments. Consequently, all DenStream-related results reported in this paper were obtained using pyDenStream, which more closely follows the original algorithmic specification and avoids repeated reclustering during prediction.
The experimental results reveal a fundamental tension between cluster quality and operational characteristics. While DenStream provides superior cluster quality metrics, the elevated Narrative Contingency and Variance suggest that incremental clustering introduces temporal instability in topic structure, a critical consideration for operational intelligence systems where analysts expect consistent topic definitions across reporting periods. It is important to note that the differences observed in narrative specific metrics may be partially attributed to the fact that the narrative extraction step (LLM-based topic labeling and enrichment) was fine tuned on the output of HDBSCAN clustering. This means the narrative metrics may inherently favor HDBSCAN’s cluster structure, as the downstream processing pipeline was optimized for the characteristics of HDBSCAN outputs. The higher Narrative Contingency and Variance observed in online algorithms may therefore reflect not only their temporal adaptation behavior but also a mismatch with the expectations encoded in the narrative extraction layer.
The baseline HDBSCAN’s strength lies in its temporal stability (low contingency and variance), achieved through complete daily re-clustering that establishes consistent reference frames. However, this stability comes at the cost of computational scalability and inability to leverage historical patterns.
DenStream emerges as the most promising online alternative, offering substantial improvements in cluster quality that may justify the trade-off in temporal stability. The algorithm’s ability to discover more distinct narrative clusters (higher distinctness) while maintaining reasonable internal coherence positions it as a viable candidate for production deployment, particularly in scenarios where cluster quality takes precedence over strict temporal consistency.
6 Conclusion and Future Work
Overall, the results indicate that density-based online clustering can serve as a viable replacement for batch HDBSCAN in narrative monitoring pipelines, provided that system design accounts for the trade-off between cluster quality and temporal stability. Among the evaluated methods, DenStream offers the most favorable balance, achieving superior cluster separation and narrative distinctness while maintaining feasible computational performance for streaming deployment. We therefore recommend DenStream as the primary candidate for production integration in systems where scalability and adaptation to evolving narratives are critical.
Further investigation should examine whether DenStream’s temporal variability can be mitigated through hyperparameter tuning or post-processing stabilization techniques, potentially enabling the algorithm to achieve both superior cluster quality and operational stability required for narrative monitoring applications. Additionally, retuning the narrative extraction pipeline to accommodate the characteristics of online clustering outputs may help reduce the observed narrative metric discrepancies and provide a more fair comparison across algorithms.
Limitations
Several limitations should be considered when interpreting the results of this study.
All experiments were conducted on data from the Ukrainian information space, characterized by high geopolitical intensity, multilingual content (Russian, Ukrainian, code-switched surjik), and elevated narrative volatility of wartime news stream. While this setting is representative of narrative monitoring applications, results may not directly generalize to domains with slower topic evolution, more formal language (e.g., scientific corpora), or lower noise levels. The relative advantages of online clustering are less pronounced in more static environments.
Clustering performance is inherently conditioned on upstream representation learning and dimensionality reduction. The experiments rely on transformer-based embeddings and UMAP projections; different encoder models, reduction techniques, or parameter settings could alter cluster geometry and affect the comparative performance of HDBSCAN and online methods. Thus, conclusions concern the clustering stage within this pipeline, not clustering in isolation.
Narrative-specific metrics are partly influenced by the LLM-based topic labeling and enrichment layer, which was originally tuned on HDBSCAN outputs. This may introduce structural bias favoring the baseline and partially penalizing online algorithms whose cluster shapes differ. Consequently, narrative metric comparisons may reflect both clustering behavior and downstream pipeline compatibility rather than pure cluster structure quality.
Some observed performance differences arise from practical limitations of available implementations (e.g., River), rather than intrinsic algorithmic properties. Although alternative implementations were used where necessary, system-level conclusions are influenced by software maturity and engineering trade-offs, which may change as libraries evolve.
The streaming simulation spans a limited historical window and a single evaluation day. While sufficient to study incremental behavior, longer-term dynamics such as gradual concept drift, seasonal topic cycles, or rare narrative shocks were not systematically evaluated. Extended longitudinal studies are required to fully assess stability and adaptation trade-offs.
Clusters are treated as proxies for narratives, yet narrative structure is inherently fluid and may not align perfectly with density-based cluster boundaries. Some differences in “narrative quality” metrics may therefore reflect modeling assumptions rather than true differences in narrative discovery capability.

### v2

@@
 Abstract
-Automated narrative intelligence systems for social media monitoring face significant scalability challenges when processing continuous data streams using traditional batch clustering algorithms.
-We investigate the replacement of HDBSCAN (offline clustering) with online (streaming/incremental) clustering methods in a production narrative report generation pipeline.
-The proposed system employs a three-stage architecture (data collection, modeling, dashboard generation) that processes thousands of multilingual social media documents daily.
-While HDBSCAN excels at discovering hierarchical density-based clusters and handling noise, its batch-only nature necessitates complete retraining for each time window, resulting in memory constraints, computational inefficiency, and inability to adapt to evolving narratives in real-time.
-This work evaluates a bunch of online clustering algorithms across dimensions of cluster quality preservation, computational efficiency, memory footprint, and integration compatibility with existing workflows.
-We propose evaluation criteria that balance traditional clustering metrics (Silhouette Coefficient, Davies-Bouldin Index) with narrative metrics (narrative distinctness, contingency and variance).
-Our methodology includes sliding-window simulations on historical datasets from Ukraine information space, enabling comparative analysis of algorithmic trade-offs in realistic operational contexts.
-This research addresses a critical gap between batch-oriented topic modeling frameworks and the streaming nature of social media monitoring, with implications for computational social science, crisis informatics, and narrative surveillance systems.
+Automated narrative intelligence systems for social media monitoring face significant scalability challenges when relying on batch clustering methods to process continuous data streams.
+We investigate replacing offline HDBSCAN with online density-based clustering algorithms in a production narrative report generation pipeline that processes large volumes of multilingual social media data.
+While HDBSCAN effectively discovers hierarchical clusters and handles noise, its batch-only nature requires full retraining for each time window, limiting scalability and real-time adaptability.
+We evaluate online clustering methods with respect to cluster quality, computational efficiency, memory footprint, and integration with downstream narrative extraction.
+Our evaluation combines standard clustering metrics, narrative-specific measures, and human validation of cluster correctness to assess both structural quality and semantic interpretability.
+Experiments using sliding-window simulations on historical data from the Ukrainian information space reveal trade-offs between temporal stability and narrative coherence, with DenStream achieving the strongest overall performance.
+These findings bridge the gap between batch-oriented clustering approaches and the streaming requirements of large-scale narrative monitoring systems.
 Online Density-Based Clustering for Real-Time Narrative Evolution Monitoring
-Ostap Vykhopen, Viktoria Skorik, Maxim Tereschenko, Veronika Solopova ostap.vykhopen@mantisanalytics.com
+Ostap Vykhopen, Viktoria Skorik, Maksym Tereshchenko, Veronika Solopova ostap.vykhopen@mantisanalytics.com
 1 Introduction
@@
 (2023).
-Despite its theoretical advantages, including the ability to handle varying density clusters, detect noise, and provide hierarchical structure, HDBSCAN’s strict batch processing requirement creates significant operational bottlenecks in production environments: (i) Memory constraints.
-Each clustering operation requires loading complete document embeddings (N × 100 dimensional UMAP reduced vectors) into memory, with O(NlogN) average case complexity for minimum spanning tree construction; (ii) computational inefficiency.
-For example, daily monitoring require complete re-clustering of new data each day, discarding previously learned cluster structures even as narrative topics evolve gradually; (iii) Parameter sensitivity.
-Production logs reveal frequent cluster size parameter tuning, indicating difficulty in maintaining stable cluster granularity across varying data volumes; (iv) inability to handle incremental updates.
-When new documents arrive, the entire clustering process must restart, preventing real-time cluster assignment or progressive refinement of narrative understanding; (v) scalability limitations.
-As historical data accumulates, the batch processing approach becomes increasingly untenable.
-These limitations motivate our central research question: Can online clustering algorithms effectively replace HDBSCAN in narrative monitoring pipelines while preserving cluster quality, maintaining system integration compatibility, and improving computational efficiency?
-Specifically, we test clustering algorithms that not only support online mode but also automatically determine the number of clusters without requiring predefined cluster counts (as required by batch K-means) and demonstrate robustness to noise inherent in social media data streams.
-Furthermore, the ideal online clustering approach should achieve significant efficiency improvements, where the memory footprint of the online algorithm is substantially smaller than HDBSCAN’s batch requirements, and the time required for incremental updates is considerably faster than HDBSCAN’s complete re-clustering operations.
+Despite its theoretical strengths, handling variable-density clusters, noise, and hierarchical structure, HDBSCAN introduces significant operational bottlenecks in production narrative monitoring.
+Its strict batch processing requires loading all document embeddings into memory for each clustering run, leading to high memory usage and repeated O(NlogN) computations.
+In daily monitoring settings, this necessitates full re-clustering of incoming data, discarding previously learned structure even when narratives evolve gradually.
+In practice, maintaining stable cluster granularity further requires frequent parameter tuning as data volume fluctuates.
+Most critically, HDBSCAN cannot support incremental updates, preventing real-time cluster assignment and progressive narrative refinement.
+As historical data accumulates, these limitations render batch re-clustering increasingly impractical at scale.
+These constraints motivate our central research question: Can online clustering algorithms replace HDBSCAN in narrative monitoring pipelines while preserving cluster quality, maintaining integration compatibility, and improving computational efficiency?
+We evaluate online density-based clustering methods that automatically infer the number of clusters, remain robust to noise in social media streams, and support efficient incremental updates with substantially lower memory and computational overhead than batch re-clustering.
 Our contributions are threefold:
   1. 1.
-    We introduce a realistic experimental framework for evaluating clustering under streaming conditions, including temporal pretraining, incremental daily updates, and sliding-window simulations on high-volume multilingual social media data.
+    We propose a realistic evaluation framework for streaming clustering, incorporating temporal pretraining, incremental daily updates, and sliding-window simulations on large-scale multilingual social media data.
   2. 2.
-    We conduct a systematic comparison of online density-based algorithms (DBSTREAM, DenStream) against the HDBSCAN baseline within a production-style pipeline, using a joint evaluation scheme that combines traditional clustering metrics (Silhouette, Davies–Bouldin) with narrative-specific metrics (distinctness, contingency, variance)
+    We systematically compare online density-based methods (DBSTREAM, DenStream) against a batch HDBSCAN baseline within a production-style pipeline, using a unified evaluation that combines standard clustering metrics with narrative-specific measures.
   3. 3.
-    We identify implementation-level scalability and correctness issues in widely used reference libraries and show how these affect real-world performance.
-Based on empirical and operational analysis, we position DenStream as the most promising online alternative and outline the architectural and downstream adaptations required for deployment in streaming narrative monitoring systems.
+    We analyze scalability and correctness issues in widely used reference implementations and demonstrate their impact on real-world performance.
+Based on empirical and operational evidence, we identify DenStream as the most viable online alternative and outline the system adaptations required for deployment in narrative monitoring pipelines.
 2 Theoretical Background and Related work
 2.1 Density-Based Clustering
-Density-based clustering algorithms, pioneered by Density-Based Spatial Clustering of Applications with Noise (DBSCAN) from Ester et al.
-(1996), have become foundational for discovering arbitrary-shaped clusters in spatial data while robustly handling noise.
-Unlike partition-based methods (k-means) or hierarchical agglomerative clustering, DBSCAN defines clusters as dense regions separated by sparse areas, requiring only two parameters: $\epsilon$ (neighborhood radius) and min_samples (minimum points to form a dense region).
-HDBSCAN extends DBSCAN by eliminating the need for $\epsilon$ through hierarchical cluster extraction based on cluster stability.
-The algorithm constructs a minimum spanning tree over mutual reachability distances, then extracts a flat clustering by selecting stable branches of the hierarchical tree.
-This approach provides several advantages: (1) automatic determination of optimal cluster granularity, (2) varying density cluster detection, (3) hierarchical cluster structure, and (4) probabilistic soft cluster assignments via outlier scores.
-Despite these strengths, traditional implementations of DBSCAN and HDBSCAN require access to the complete dataset to construct their distance-based structures and compute cluster hierarchies.
-This computational model assumes static data and the ability to hold all points in memory simultaneously, limiting applicability to streaming scenarios where data arrives continuously and retrospective reprocessing of the entire history is impractical.
-Furthermore, narrative monitoring applications, where topics emerge, evolve, and fade over time, demand algorithms that can discover clusters without predefining their number, adapting dynamically to the natural structure present in incoming document streams rather than imposing predetermined partitions.
+Density-based clustering, introduced with DBSCAN (Ester et al., 1996), identifies clusters as dense regions separated by sparse areas, enabling the discovery of arbitrarily shaped clusters while robustly handling noise.
+Unlike partition-based methods such as k-means or hierarchical agglomerative clustering, DBSCAN requires only two parameters—the neighborhood radius $\epsilon$ and a minimum density threshold—and does not require predefining the number of clusters.
+HDBSCAN extends DBSCAN by removing the need to specify $\epsilon$ through hierarchical clustering based on mutual reachability distances and cluster stability.
+By extracting stable branches from a minimum spanning tree, HDBSCAN automatically determines cluster granularity, supports varying-density clusters, and provides soft cluster assignments via outlier scores.
+Despite these advantages, DBSCAN and HDBSCAN are inherently batch-oriented, requiring access to the full dataset to construct distance structures and extract clusters.
+This assumption limits their applicability to streaming settings, where data arrives continuously and full reprocessing is impractical.
+In narrative monitoring applications, where topics emerge, evolve, and fade over time, clustering methods must adapt incrementally without predefined cluster counts, motivating the use of online density-based alternatives.
 2.2 Online and Streaming Clustering Algorithms
-The streaming clustering literature addresses scenarios where data arrives continuously and memory constraints prevent storing the entire dataset.
-CluStream Chan et al.
+2.3 Streaming Clustering
+Streaming clustering addresses scenarios in which data arrives continuously and memory constraints preclude storing the full dataset.
+Early work such as CluStream Chan et al.
 (1982); Caussinus et al.
-(1982) is one of the earliest and most influential data stream clustering frameworks, introducing a two-phase approach consisting of online micro-clustering and offline macro-clustering.
-In the online phase, it maintains temporal micro-clusters using summary statistics and timestamps, while the offline phase applies traditional clustering algorithms (e.g., k-means) over these summaries to generate final clusters at different time horizons.
-While CluStream supports temporal analysis and evolving streams, it requires the number of clusters to be specified in advance and is primarily suited for spherical cluster structures.
-DBSTREAM Hahsler and Bolanos (2016) is a density-based data stream clustering algorithm that extends micro-cluster approaches by maintaining a shared density graph capturing the actual density between adjacent micro-clusters during the online summarization.
-At reclustering time, it uses this shared density information to connect micro-clusters into final clusters more accurately than methods relying only on distances or assumptions about internal distributions.
+(1982) introduced a two-phase framework combining online micro-clustering with offline macro-clustering.
+While CluStream enables temporal analysis of evolving streams, it requires the number of clusters to be specified in advance and is best suited to spherical cluster structures.
+Subsequent density-based approaches remove this limitation.
+DBSTREAM Hahsler and Bolanos (2016) maintains a shared density graph between micro-clusters, allowing final clusters to be reconstructed based on density connectivity rather than distance alone.
 DenStream Cao et al.
-(2006) is another density-based stream clustering maintaining micro-clusters (core-micro-clusters and outlier-micro-clusters) with temporal decay weights.
-Its crucial advantage is that it addresses concept drift by gradually forgetting old data and it supports arbitrary cluster shapes like DBSCAN.
-However, this introduces additional complexity in managing micro-cluster lifecycle and decay parameters.
-Recent neural approaches, including deep clustering networks and variational autoencoders with clustering objectives, offer representation learning combined with clustering, yet they typically require large training datasets and provide limited interpretability, which is essential for narrative monitoring applications.
+(2006) extends this idea by incorporating temporal decay through core and outlier micro-clusters, enabling adaptation to concept drift while supporting arbitrary cluster shapes.
+However, this flexibility introduces additional complexity in managing micro-cluster lifecycles and decay parameters.
+More recent neural clustering approaches combine representation learning with clustering objectives, but they typically require large training datasets and offer limited interpretability, which constrains their applicability in analyst-facing narrative monitoring systems.
 2.4 Streaming Topic Modeling
-Traditional topic modeling methods such as LDA and NMF have been extended to streaming and online contexts using incremental variational inference and related techniques, and researchers have also developed enhanced versions that combine topic models with semantic representations Das (2025).
-However, the canonical implementations of LDA and NMF still operate on bag‑of‑words or weighted document‑term matrices and do not inherently capture deep contextual similarity like transformer‑based dense embeddings do, motivating hybrid models that merge topic modeling with embedding spaces for richer semantic coherence Yadav et al.
+Traditional topic models such as LDA and NMF have been extended to online settings through incremental inference and hybrid approaches Das (2025).
+However, these methods primarily rely on bag-of-words representations and do not naturally capture contextual semantic similarity.
+Recent work therefore integrates topic modeling with dense embeddings to improve semantic coherence Yadav et al.
 (2025).
-Recent advances in neural embedding models have enabled a fundamentally different approach to streaming topic modeling that leverages the rich semantic representations of transformer-based architectures Angelov (2020); Bianchi et al.
+An alternative embedding-first paradigm leverages transformer-based representations for streaming topic discovery Angelov (2020); Bianchi et al.
 (2021).
-Rather than relying on term-frequency statistics, our approach builds on a pipeline that transforms documents into dense embeddings, applies dimensionality reduction to preserve local semantic structure in a lower-dimensional space, performs clustering to identify topically coherent groups, and finally employs large language models to generate interpretable topic labels and summaries.
-This embedding-first paradigm naturally captures contextual similarity and semantic nuance while remaining compatible with streaming scenarios through incremental clustering updates.
-By decoupling representation learning from cluster assignment, this framework addresses key limitations of traditional topic models, such as their inability to capture deep semantic relationships, while maintaining the flexibility needed for online, continuously arriving data streams.
+In this framework, documents are embedded into a dense semantic space, clustered to identify coherent topics, and subsequently labeled or summarized using language models.
+By decoupling representation learning from cluster assignment, this approach captures deep semantic relationships while remaining compatible with incremental clustering, making it well suited for large-scale narrative monitoring over continuously arriving data streams.
 3 Methodology & Experimental setup
@@
 We evaluate online clustering algorithms using data from the Ukrainian information space, including news portals, Telegram channels, and user comments.
-The dataset [^2] The data can be shared upon request to vetted academics for reproducibility purposes.
+The dataset [^2] The data and code can be shared upon request to vetted academics for reproducibility purposes.
 exhibits typical characteristics of narrative monitoring applications: high volume ($\approx$17,000 documents per day), topical diversity reflecting ongoing geopolitical discourse, and substantial noise from low-quality or off-topic content.
@@
 DenStream 0.685 0.453 0.319 0.389 0.319
-Table 2: Operational Metrics.
-The number of narrative candidates does not represent the actual number of narratives, but rather the number of clusters classified as narratives by the model.
-This distinction should be considered when interpreting the results.
 Algorithm Train (s) Pred. (s) Clusters Narratives
@@
 DenStream 3.56 1.7 303 53
+Table 2: Operational Metrics.
+The number of narrative candidates does not represent the actual number of narratives, but rather the number of clusters classified as narratives by the model.
+This distinction should be considered when interpreting the results.
 (a) HDBSCAN
@@
 Conversely, DenStream may tend to merge several distinct narratives into a single cluster.
-5 Discussion
-Overall, our results indicate that DenStream provides the most robust clustering performance, achieving clear improvements over the HDBSCAN baseline in both conventional and narrative-specific metrics.
+4.3 Human Evaluation
+Method $\overline{\kappa}$ $\alpha$ Accepted Accuracy
+HDBSCAN 0.731 0.733 183 / 236 0.775
+DBStream 0.784 0.781 94 / 116 0.810
+DenStream 0.827 0.829 89 / 106 0.840
+Table 3: Human evaluation of narrative clustering quality.
+Three annotators judged whether each generated narrative cluster was correct (true) or incorrect (false).
+We report mean pairwise Cohen’s $\kappa$ ($\overline{\kappa}$), Krippendorff’s $\alpha$ (nominal), and consensus-based acceptance using majority vote, and accuracy of the clustering based on acceptance rate.
+Table 3 summarizes the human evaluation of narrative clustering quality across three clustering methods.
+Inter-annotator agreement is consistently high, with mean pairwise Cohen’s $\overline{\kappa}$ ranging from 0.73 to 0.83 and Krippendorff’s $\alpha$ between 0.73 and 0.83, indicating substantial to strong agreement.
+Using majority vote as a proxy for cluster acceptance, DenStream achieves the highest accuracy (0.84), followed by DBStream (0.81) and HDBSCAN (0.78).
+These results indicate that DenStream produces narrative clusters that are more frequently judged as coherent and correct by human evaluators.
+Notably, methods with higher consensus acceptance rates also exhibit higher inter-annotator agreement.
+In particular, DenStream achieves both the highest acceptance rate and the strongest agreement($\overline{\kappa}=0.83$, $\alpha=0.83$), suggesting that its narrative clusters are not only more frequently judged as correct, but also less ambiguous and more consistently interpretable by human evaluators.
+5 Discussion and Conclusion
+Overall, our results indicate that DenStream provides the most robust clustering performance, achieving clear improvements over the HDBSCAN baseline across both automatic clustering metrics and human evaluation.
 Its higher Silhouette Score and lower Davies–Bouldin Index suggest well-separated and cohesive clusters, which is further reflected in its superior Narrative Distinctness, highlighting its ability to recover meaningful and distinct thematic structures.
+Crucially, these structural advantages are corroborated by human evaluation: DenStream achieves the highest consensus acceptance rate (0.84) and the strongest inter-annotator agreement, indicating that its narrative clusters are not only quantitatively well-formed but also consistently interpretable and semantically coherent to human evaluators.
 In contrast, DBStream performs poorly across standard clustering metrics, likely due to limitations of its micro-cluster representation in high-dimensional semantic embedding spaces.
-While the baseline exhibits greater temporal stability, as indicated by lower Narrative Contingency and Variance, the increased variability observed for the online algorithms appears to be an inherent trade-off of incremental clustering.
-This adaptability may be advantageous in dynamic intelligence settings, where evolving narratives are expected rather than static topic structures.
-However, during the implementation and experimental evaluation of clustering algorithms, several technical limitations were identified in commonly used reference implementations, substantially affecting performance and scalability, which should be taken into consideration when choosing a method.
-For DenStream, the predict_one method in the River ML library recomputes final clusters by executing a DBSCAN-like procedure over all micro-clusters on every prediction call.
-While this behavior is consistent with fully online streaming assumptions, it leads to prohibitively slow prediction when the model state is static.
-This limitation is acknowledged by the River maintainers, who note that ‘DenStream.predict_one‘ performs excessive recomputation for certain use cases.
-In offline evaluation settings, this repeated reclustering is unnecessary.
-Additionally, inconsistencies were observed in River’s DenstreamMicroCluster implementation, particularly in the maintenance of running weighted statistics, which can result in incorrect micro-cluster centers, radii, and weights.
-For CluStream, performance issues arise during training rather than prediction.
-When micro-cluster merging is required, the algorithm performs quadratic-time pairwise distance computations between all micro-cluster centers, resulting in (O($M^{2}$)) complexity per merge event Mansalis et al.
-(2018).
-Combined with additional overhead from periodic k-means reclustering, this leads to rapidly increasing training time as the stream progresses.
-Due to these inherent scalability limitations, CluStream was fully rejected from further consideration in this study.
-As a result of the identified performance and correctness issues, the River implementations were deemed unsuitable for large-scale experiments.
-Consequently, all DenStream-related results reported in this paper were obtained using pyDenStream, which more closely follows the original algorithmic specification and avoids repeated reclustering during prediction.
-The experimental results reveal a fundamental tension between cluster quality and operational characteristics.
-While DenStream provides superior cluster quality metrics, the elevated Narrative Contingency and Variance suggest that incremental clustering introduces temporal instability in topic structure, a critical consideration for operational intelligence systems where analysts expect consistent topic definitions across reporting periods.
-It is important to note that the differences observed in narrative specific metrics may be partially attributed to the fact that the narrative extraction step (LLM-based topic labeling and enrichment) was fine tuned on the output of HDBSCAN clustering.
-This means the narrative metrics may inherently favor HDBSCAN’s cluster structure, as the downstream processing pipeline was optimized for the characteristics of HDBSCAN outputs.
-The higher Narrative Contingency and Variance observed in online algorithms may therefore reflect not only their temporal adaptation behavior but also a mismatch with the expectations encoded in the narrative extraction layer.
-The baseline HDBSCAN’s strength lies in its temporal stability (low contingency and variance), achieved through complete daily re-clustering that establishes consistent reference frames.
-However, this stability comes at the cost of computational scalability and inability to leverage historical patterns.
-DenStream emerges as the most promising online alternative, offering substantial improvements in cluster quality that may justify the trade-off in temporal stability.
-The algorithm’s ability to discover more distinct narrative clusters (higher distinctness) while maintaining reasonable internal coherence positions it as a viable candidate for production deployment, particularly in scenarios where cluster quality takes precedence over strict temporal consistency.
-6 Conclusion and Future Work
-Overall, the results indicate that density-based online clustering can serve as a viable replacement for batch HDBSCAN in narrative monitoring pipelines, provided that system design accounts for the trade-off between cluster quality and temporal stability.
-Among the evaluated methods, DenStream offers the most favorable balance, achieving superior cluster separation and narrative distinctness while maintaining feasible computational performance for streaming deployment.
-We therefore recommend DenStream as the primary candidate for production integration in systems where scalability and adaptation to evolving narratives are critical.
-Further investigation should examine whether DenStream’s temporal variability can be mitigated through hyperparameter tuning or post-processing stabilization techniques, potentially enabling the algorithm to achieve both superior cluster quality and operational stability required for narrative monitoring applications.
-Additionally, retuning the narrative extraction pipeline to accommodate the characteristics of online clustering outputs may help reduce the observed narrative metric discrepancies and provide a more fair comparison across algorithms.
+While DBStream achieves a moderate acceptance rate in human evaluation, its lower mean pairwise Cohen’s $\kappa$ compared to DenStream suggests that its cluster outputs are more ambiguous, making it harder for annotators to consistently agree on their correctness.
+This highlights an important distinction between acceptance frequency and interpretability: clusters may be judged acceptable in isolation while still lacking the clarity that enables stable human agreement.
+The baseline HDBSCAN exhibits lower cluster distinctness in both automatic metrics and human acceptance rate, yet demonstrates greater temporal stability, as indicated by lower Narrative Contingency and Variance.
+This stability arises from complete daily re-clustering, which enforces consistent reference frames across time.
+However, human evaluation reveals that this stability does not necessarily translate into higher perceived cluster quality, suggesting that static re-clustering may preserve structural consistency at the expense of semantic sharpness.
+Taken together, the experimental results reveal a fundamental tension between cluster quality, interpretability, and temporal stability.
+While DenStream introduces higher Narrative Contingency and Variance due to its incremental nature, human evaluation indicates that this variability corresponds to improved semantic clarity rather than noise.
+In dynamic intelligence settings, where evolving narratives are expected, such adaptability may be preferable to enforcing static topic definitions.
+Finally, it is important to note that narrative-specific metrics may partially reflect biases introduced by the narrative extraction pipeline, which was fine-tuned on HDBSCAN outputs.
+As a result, some discrepancies observed for online algorithms may stem from a mismatch between clustering behavior and downstream expectations rather than intrinsic deficiencies in cluster quality.
+The stronger human validation of DenStream clusters suggests that these narrative metrics underestimate the semantic validity of online clustering approaches.
 Limitations
@@
 Some differences in “narrative quality” metrics may therefore reflect modeling assumptions rather than true differences in narrative discovery capability.
+DenStream’s temporal variability might be mitigated through hyperparameter tuning or post-processing stabilization techniques, potentially enabling the algorithm to achieve both superior cluster quality and operational stability required for narrative monitoring applications.
+Additionally, retuning the narrative extraction pipeline to accommodate the characteristics of online clustering outputs may help reduce the observed narrative metric discrepancies and provide a more fair comparison across algorithms.
+Appendix A Implementation Challenges
+During the implementation and experimental evaluation of clustering algorithms, several technical limitations were identified in commonly used reference implementations, substantially affecting performance and scalability, which should be taken into consideration when choosing a method.
+For DenStream, the predict_one method in the River ML library recomputes final clusters by executing a DBSCAN-like procedure over all micro-clusters on every prediction call.
+While this behavior is consistent with fully online streaming assumptions, it leads to prohibitively slow prediction when the model state is static.
+This limitation is acknowledged by the River maintainers, who note that DenStream.predict_one performs excessive recomputation for certain use cases.
+In offline evaluation settings, this repeated reclustering is unnecessary.
+Additionally, inconsistencies were observed in River’s DenstreamMicroCluster implementation, particularly in the maintenance of running weighted statistics, which can result in incorrect micro-cluster centers, radii, and weights.
+For CluStream, performance issues arise during training rather than prediction.
+When micro-cluster merging is required, the algorithm performs quadratic-time pairwise distance computations between all micro-cluster centers, resulting in $O(M^{2})$ complexity per merge event Mansalis et al.
+(2018).
+Combined with additional overhead from periodic k-means reclustering, this leads to rapidly increasing training time as the stream progresses.
+Due to these inherent scalability limitations, CluStream was fully rejected from further consideration in this study.
+As a result of the identified performance and correctness issues, the River implementations were deemed unsuitable for large-scale experiments.
+Consequently, all DenStream-related results reported in this paper were obtained using pyDenStream, which more closely follows the original algorithmic specification and avoids repeated reclustering during prediction.